OpenStudy (safia21):

A riding lawnmower can drive up a 27º incline. In order for a ramp designed to reach a truck bed that is 3 feet above the ground to have a 27º incline, how long should the ramp be?

OpenStudy (anonymous):

think of it as a right triangle

OpenStudy (anonymous):

and use tanx

OpenStudy (anonymous):

think trigonometry if ABC a right triangle with B the right angle, then A should be 27 degrees, and side BC should be a vertical line from the truck to the earth equal to 3 feet. The other angle will be 63o. Use sins and cosins

OpenStudy (anonymous):

3/x = tan(27o)

OpenStudy (anonymous):

The tan of an angle is defined as the ratio of the length opposite the angle (27 degrees in this case) to the length of the hypotenuse of the right triangle (the length of the ramp, Wallach used x). Thus, \[\tan27^{o}=x/3\] or \[x=3\tan27^o \approx1.5286\]

OpenStudy (anonymous):

mandolino im sorry my math terminology in english is not that good but i remember that tan is the side of the triangle that is in front of the angle divided by the side of the triangle which is to the side of the angle which will make it 3/x and not x/3

OpenStudy (anonymous):

3 feet ABOVE

OpenStudy (anonymous):

You are correct. Thank you. The correct should be: The tan of an angle is defined as the ratio of the length opposite the angle (27 degrees in this case) to the length of the hypotenuse of the right triangle (the length of the ramp, Wallach used x). Thus, \[\tan27^o=3/x\] or \[x=3/\tan27^o \approx5.888\]